OpenStudy (anonymous):

l5-4xl-3=4 (Solving Absolute-Value Equations and Inequalities)

OpenStudy (dumbcow):

\[|5-4x|-3=4\] \[|5-4x| = 7\] \[-7 = 5-4x = 7\] \[-12 = -4x = 2\] \[3 = x = -1/2\] 2solutions: x=3, x = -1/2